A monthly salary is $1371. It is given an increase of 8.3%. After the first increase, it receives a second increase of 10%. What is the new monthly salary after both increases? Round to the nearest cent.

Answer by Theo(13342)

the monthly salary is 1371.
to apply an increase of 8.3%, multiply by 1.083.
to apply an increase of 10%, multiply by 1.10.
to apply both increases multiply by 1.08 * 1.10
you will get 1371 * 1.08 * 1.10 = 1633.2723
your first increase got you from 1371 to 1371 * 1.083 = 1484.793.
your second increase got you from 1484.793 to 1484.793 * 1.10 = 1633.2723.